OpenStudy (anonymous):

Find the integral of x(6x+7)^8 dx

OpenStudy (raden):

integral by subs, let u = 6x+7

OpenStudy (anonymous):

Yes I tried that but nothing canceled out

OpenStudy (anonymous):

because du/dx would equal to 6

OpenStudy (raden):

u = 6x+7 ---------> x=(u-7)/6, right ? du=6dx or dx = 1/6 * du now, plug of them to the original problem

OpenStudy (anonymous):

oh I see. I got dx=6du instead of 1/6 du

OpenStudy (raden):

btw, u were missing dx above :)

OpenStudy (anonymous):

haha thanks, i'll add it in

OpenStudy (raden):

welcome

OpenStudy (anonymous):

so now I got integral of (u-7)(u^8) du. Is that right?

OpenStudy (raden):

where's 1/6 ?

OpenStudy (anonymous):

oops I meant integral of [ (u-7)/36 ] (u^8) du

OpenStudy (anonymous):

so its 1/36 integral of (u-7)(u^8) du

OpenStudy (raden):

ok...

OpenStudy (anonymous):

Still can't integrate

OpenStudy (raden):

use distributive property to simplies (u-7)u^8 = u^9 - 7u^8

OpenStudy (raden):

just integrate it one by one

OpenStudy (anonymous):

\[\int\limits x(6x+7)^8 dx\]\[u=6x+7\]\[{du \over 6}=dx\]\[x={u-7 \over 6}\]\[\int\limits {u-7 \over 6}u^8{du \over 6}\]\[\int\limits {u^9-7u^8 \over 36}du\]\[{1 \over 36}\int\limits u^9-7u^8 \space du\]I'll let you finish

OpenStudy (anonymous):

Thanks got my answer. [ (6x+7)^10 /360 ] - [ 7(6x+7)^9 /324]

OpenStudy (raden):

yeah, correct

OpenStudy (raden):

dont forget + c :p

OpenStudy (anonymous):

This was a tricky one.. :o thanks for sticking by lol